OpenStudy (anonymous):

find f o g f(x)=x+11 g(x)=2+x^2

OpenStudy (triciaal):

use the expression for g(x) as x in the function f(x) f (x) = x + 11 f o g = 2+x^2 + 11 then simplify if possible